MARS DESERT RESEARCH STATION

GreenHab Report – December 22nd

GreenHab Report:

Today marked the completion of the transplantation effort in the GreenHab.  We have assembled eight total conveyor trays (four Red Oak Lettuce and four Green Oak Lettuce in both hydroponic solution and in soil) and one tray with a density experiment which was planted in soil.  The slow process of sealing the GreenHab also began today.  We have continued to notice that despite the near continuous operation of the heater, the GreenHab does not stay within an acceptable temperature range.  At night, the temperature gets down to between 48 F and 54 F which is not conducive to productive and rapid plant growth that is necessary to sustain a healthy astronaut’s diet.  In order to help prevent some of the theorized leakage of heated air, clear caulk is being used to seal the places at the top of the GreenHab where air could escape.  This is a difficult process as the materials from which the GreenHab is constructed are clearly not meant to be used for a greenhouse.  We are concerned there may not be enough caulk on Mars to seal the GreenHab.  I hope future crews can figure out how to stabilize the temperature. Finally, the seedlings are germinating at excellent rates!  We can easily identify growth in the radish, green and red oak lettuce, carrot, and mystery crop.  Yay mystery crop!​

Thank you,

-Sean Gellenbeck

SEDS-MDRS Crew HSO

 SEDS-MDRS Crew GreenHab Officer #2

Journalist Report – December 22nd

Sol 04
Journalist report to be posted
Authored by Anselm Wiercioch

It snowed for a few hours today. That’s not supposed to happen.

Going to have to put some serious effort into revamping our
understanding of Mars’ climate. Coldest day since we landed. The EVA
crew decided to postpone until tomorrow due to potentially inclement
weather. I’m not on tomorrow’s crew, but we might need to reprioritize
some climate data while they’re out.. Will keep investigating.

Otherwise, today was a slow day. Some progress was made in the
greenhouse, and the network connectivity issues persist despite many
hours of messing with it. Felt like a snow day in elementary school
where you’re off class and can feel christmas around the corner.

Our crew engineer engineered some cinnamon raisin swirl bread and it’s
magical. The freeze dried food stores will probably start to wear on
us eventually, but for the time being we’re living it up. As long as
coffee and tea holds out, crew morale is going to be coasting just
fine.

Connor and I’s sleep schedule is holding out strong. We aren’t being
too strinct about the schedule and we aren’t going too extreme – still
a solid six hours or so per night, spaced into ~3 naps. We’ve been
ever so slightly tired but that’s to be expected on the first day or
two. Generally feel pretty energized though.

The most annoying thing at the moment for me is just being cooped up
in a tin can. Really struggling to understand how all those super
smart engineers on the ground decided a treadmill wasn’t necessary.
Yoga and pushups only get you so far. The hab’s air system isn’t
exactly refreshing either. Meh. It’s all good though. I’m sure we’ll
get used to it. Or at least, we’ll be gone before it really starts to
get to us.

We did find a massive binder of awesome (mostly old and/or super
goofy) movies, so that’s helping the nights pass faster after work is
done each day.

Nothing too crazy. The days are moving by faster as a whole.

#Snowday

Sol Summary – December 22nd

Daily Summary Report
SOL: 04
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Snowday
Mission Status: Active. Full crew functional.
Sol Activity Summary: Snow on ground, greenhouse work continues, EVA postponed
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: Nothing
Weather: High 35F, Low 29F, wind avg 1.5mpg, gust N/A, humidity
35-98%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Postponed until tomorrow
Reports to be filed:
– Sol Summary
– Journalist’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested
– None

Science Report – December 21st

Mars Self Sleep Report Study/Crew Well being:

Last night I tried to get up at 2 am but was not able to do work for
long before going back to sleep. I know our plan was to only sleep for
three hours at night and nap during the day but we found that the
first few days of adjusting to this process are the worst. Almost as
bad as flying to Europe or Asia through may time zones and adjusting.

We want to raise awareness of the time differences for a Mars mission
compared to living on Earth. How will the greater than 24 hr day on
Mars translate to astronauts work and sleep patterns? Obviously on
Mars, astronauts will be working and fixing things every minute while
they are awake. We want to see how we can maximize astronauts’
productivity.

As a crew we are going to try and go to sleep earlier one night and
start our day by 6 am instead of 8 am like we usually do to see how
this improves or reduces productivity.

The crew seems to be in great spirits doing their work and surviving
here on Mars. Cannot wait to see what SOL 4-13 bring!!

Crew Photos – December 21st

The EVA team departs
Sean inspects plants as darkness falls in the GreenHab
GreenHab officers Brittany and Sean tend to some lettuce plants
Brittany waters some new transplants
Connor takes a selfie as the EVA crew prepares to leave
Anselm prepares to release Bernie the pack rat into the
wild. Be free, Bernie!!
Connor and Alison get into the spirit of the EVA
Connor, Alison, and Anselm pose during the EVA

Commander Report – December 21st

Alison Gibson
Commander Report
21 Dec 2016 – Sol 3

Mission Support,

We’ve been on Mars for 3 sols now, and are forming a strong crew
dynamic. Everyone is doing their best to ensure a productive and
successful mission. Three of us went on an EVA today (Connor, Anselm,
and myself) out to the Dinosaur Quarry to explore the area and get
comfortable on the rovers while suited up. It’s been a cold and cloudy
day, so the ride out and back was frigid. On our way out to the
quarry, we set our pack rat (Bernie) free in the middle of the Martian
desert. After returning, we all ate lunch together and discussed our
plans for baking holiday treats this weekend.

We are generally staying warm and hydrated; although, while the
facility issues have been addressed, we are disappointed with the
internet access. Many of us were relying on internet access to
complete tasks during our mission, and I personally need wi-fi to
complete my research project. We cannot receive email or even navigate
to a page on a browser about 90% of the time. This has made evening
communications with CapComm, Mission Support, and Shannon Rupert very
difficult. We also cannot track our package of mission patches or
flight suits without the internet. Due to access settings put in place
by the previous Crew Commander, we cannot utilize the other two
routers at all. We’ve been monitoring our bandwidth, have turned off
all Bluetooth devices, taken turns using our computers, and reset the
router numerous times. Nothing has helped. This issue is causing
seriously impediments to our mission goals, and we hope there is a way
to resolve it with the help of Mission Support.

A Sol Summary, Engineering Report, HSO Report, Science Report, and
photos will follow. There is also an EVA Request for Sol 4.

Thank you for the weather report and your support.

Respectfully submitted,

Alison Gibson
Commander, MDRS Crew 171

Sol Summary – December 21st

SOL: 03
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Longer EVA, beginning greenhab setup
Mission Status: Active. Full crew functional.
Sol Activity Summary: EVA’d to Dinosaur Quarry, set up first conveyor
stage in greenhab, planted first 12 plants.
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: back side of helmet cracked slightly on rock.
Repaired with superglue and zip ties. Not majorly concerning, but
helmet henceforth moved to reserve use only.
Weather: High 42F, Low 18F,  wind avg 3.5mph, gust N/A, humidity
17-41%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Anselm, Geoff, and Alison to Dinosaur Quarry
Reports to be filed:
– Sol Summary
– Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested:
– Wifi still intermittent.

GreenHab Report – December 21st

Today began the transplantation of lettuce plants into the Greenhab!
We chose our heartiest crop as the test cultivar for GreenHab
conditions.  Temperatures have seemed to normalize but today was
completely cloud-covered and therefore hostile conditions may return
with the sun.  We are leveraging differences in soil and hydroponic
systems in order to compare which growing systems may have advantages
in an analog Martian habitat. Conveyor plant stages were setup for Red
Oak Lettuce today.  Eight developmental stages will be present by the
end of the mission. Six were carefully separated from the Rockwool in
order to minimize root damage and planted into equal volumes of wet
soil.  The hydroponic setup involved creating a hydroponic solution
suitable for the broad range of plant life stages present in the
conveyor experiment.  This involved starting with tap water and first
adding a nutrient solution to get the electrical conductivity (EC) to
an appropriate level.  This brought the EC to about 1.9 mS/cm.  The
next step was to lower the pH to a slightly basic value.  After
calibrating the sensor, the pH was reading 6.06 which was well within
the acceptable range.  If all goes well, Green Oak Lettuce and other
cultivars will be transplanted tomorrow.

Science Report – December 21st

Geology :
Today on our EVA we observed many interesting rock outcroppings and
geologic formations. It is easy to see the stratified rock layers in
exposed hillsides and cliffs. I moved the main weather station here at
the HAB from the roof to ground level outside the HAB. The reason for
this was that on the roof some of the heat from inside was affecting
the temperature measurements. On the next EVA that I go on I plan to
put another camera in a location of geologic interest (to be
determined) as well as collect the video data from the time-lapse
camera that has been sitting near the Hab for a couple days now. This
camera will have daytime geology data as well as star exposure.
Tomorrow it may snow here so I may give a camera to tomorrow’s EVA
team to set up outside somewhere.

I plan to move the weather station inside the GreenHAB within the next
couple days so that I can gather accurate weather data first. We want
to eventually move it into the GreenHAB in order to use the solar
sensor to measure solar flux in the Green HAB compared to outside.

The last time-lapse camera still sits inside the GreenHAB to monitor
progress there.

Max/Min: Outdoor Temp – 18 F – 42 F
GreenHab Temp – 47 F – 73 F
Barometer – 29.81 – 29.87 inHg
Wind – 3.5 mph, gust – N/A mph
Solar Rad. Max – 201.7 W/m^2
UV Index – 576 uW/cm^2
Outdoor Humidity – 17% – 41%

Crew Photos – December 20th

Bernie the Pack Rat

 

Brittany helps Alison Suit Up

 

Brittany hold her crew members’ shadows on Lower Blue Hill

 

EVA Team 2

 

Jumping for Joy

 

Overlooking the Hab

 

Anselm and Alison Adventure

 

EVA Team 1